OpenStudy (anonymous):

How do I solve: 4y^2 - 11y - 3 =0 by factoring?

OpenStudy (anonymous):

Well, do you know how to factor by decomposition?

OpenStudy (anonymous):

by decomposition?

OpenStudy (anonymous):

Yeah, we compose the middle term into smaller terms

OpenStudy (anonymous):

decompose*

OpenStudy (anonymous):

yes I do

OpenStudy (anonymous):

but if I do that: (2y-...)(2y-...)

OpenStudy (anonymous):

So what two numbers sum up to -11, and are multiplied to get -12?

OpenStudy (anonymous):

to get -12?

OpenStudy (anonymous):

Yeah?

OpenStudy (anonymous):

ohhh, 4 x -3

OpenStudy (anonymous):

i get it now

OpenStudy (anonymous):

Its 12, and minus one, it has to also add up to -11

OpenStudy (anonymous):

yup! okay, thank you!

OpenStudy (anonymous):

Your welcome :), ill be here for a bit if you need anymore help

OpenStudy (anonymous):

You can factor by grouping... \[4y^2-11y-3=0\] \[4y^2-12y+y-3=0\] Remove a 3 from the first 2 terms. \[4y(y-3)+(y-3)=0\] Factor out a (y-3) \[(y-3)(4y+1)=0\] Easy to find the zeroes from here.

OpenStudy (anonymous):

Remove a 4y from the first 2 terms***

OpenStudy (anonymous):

Why do I have to remove a 4y from the first 2 terms? And why is it now (y-3)(4y + 1)?